OpenStudy (anonymous):

6x^3 How do I find three factorizations.

OpenStudy (anonymous):

What exactly is your question?

OpenStudy (anonymous):

Not sure how to do this. the book tells me to find three factorizations for 6x^3

OpenStudy (anonymous):

Can you upload the page where this question is?

OpenStudy (anonymous):

May be you write it as: \[6x^3 = 2 \times 3 \times x^3\] So, the 3 factors may be 2, 3 and x..

OpenStudy (anonymous):

Er well to be more specific it says "For questions 1-9 find three factorizations for each monomial." The answer key says that answers may vary.

OpenStudy (anonymous):

@waterineyes is right and you can also right x^3 as X * X^2 or some other ways to get three factors so answers may vary

OpenStudy (anonymous):

How do I find the factorizations though? do I divide or..?

OpenStudy (anonymous):

Yep, Meinme is also right, the answer may vary according to what you want to choose as factors.. According to you , what is factor?

OpenStudy (anonymous):

See, for example, if you have 12, and I say to find factors of it then what will you do??

OpenStudy (anonymous):

Now you know that : \[12 = 3 \times 4 = 3 \times 2 \times 2\] So, 2, 3 as well as 4 are the factors of 12..

OpenStudy (skullpatrol):

@JamesR4494 any questions?